OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <8, 7>, v = <9, 7>

OpenStudy (sidsiddhartha):

use dot product \[\Large u.v=|u|*|v|*\cos \theta\]

OpenStudy (anonymous):

First, find \(a\cdot b\) using the method I just showed you in your previous question, then use the formula \[a\cdot b = |a||b|\cos\theta\], where |a| = \(\sqrt{(a_1)^2+(a_2)^2}\), and similar for |b|

OpenStudy (sidsiddhartha):

u and v can be expressed as \[u=8 \hat i+7 \hat j \\v=9 \hat i+7 \hat j\]

OpenStudy (sidsiddhartha):

now can u find \[u.v=?\]

OpenStudy (anonymous):

the lines by the letters, what do they mean?

OpenStudy (sidsiddhartha):

\[<a,b,c> they ~can~be~expressed~as \\a \hat i +b \hat j+c \hat k\] like this

OpenStudy (sidsiddhartha):

or u can keep them as they are now can u find out first (u.v)

OpenStudy (sidsiddhartha):

\[u.v=<8,7>*<9,7>=8*9+7*7=72+49=121 \\ok???\]

OpenStudy (anonymous):

i knew that but that is not what im looking for

OpenStudy (sidsiddhartha):

so which part ure not getting?

OpenStudy (sidsiddhartha):

now you just need to evaluate ,|u| and |v| can u do that?